A certain computer design has 64 different sequences of 10 binary digits so that the total number of possible states is (2^10)^64 = 1024^64. Evaluate 1024^64 using logarithms. Show your work.
Answer by stanbon(57962)

Evaluate 1024^64 using logarithms
----
Let x = 1024^64
logx = 64*log(1024)
logx = 64*3.0103
---
logx = 192.66
---
x = 10^192.66
x = overflow on my TI-84
==============================
Cheers,
Stan H.